OpenStudy (anonymous):

Find the points on the graph of y=secx, 0

OpenStudy (anonymous):

aka slope is \(\frac{2}{3}\)

OpenStudy (anonymous):

i guess you have to solve \[\sec(x)\tan(x)=\frac{2}{3}\] but it is not clear to me how to do this easily

OpenStudy (anonymous):

i bet \(\frac{\pi}{6}\) works, try it out